OpenStudy (anonymous):

HELP PLEASE D: I never understood the explicit and recursive formula. I'm suppose to make the pattern 1, 2x, 6x^2, 24^3,... into an explicit and recursive formula

OpenStudy (anonymous):

another way to look at this pattern is 1x^0, 2x^1 etc recursive is defined by a1 any number ^0 is 1 so 1x^0 =1*1=1 can you figure out recursive from there?